OpenStudy (anonymous):

The ratio of the sides EBD of to the sides of ABC is 3:4. Calculate the side lengths of ABC (Round to the nearest hundredth)

jimthompson5910 (jim_thompson5910):

BE/AB = 3/4 5/x = 3/4 5*4 = 3x 3x= 20 x = 20/3 x = 6.67 So AB = 6.67 Use this idea to find the lengths of BC and AC

OpenStudy (anonymous):

How do I do that though. Would it be BC and DC

jimthompson5910 (jim_thompson5910):

BD/BC = 3/4 4/y =3/4 solve for y to find BC

OpenStudy (anonymous):

5.333

jimthompson5910 (jim_thompson5910):

good, so BC = 5.33

jimthompson5910 (jim_thompson5910):

finally DE/AC = 3/4 4.5/z = 3/4 solve for z, Round to the nearest hundredth

OpenStudy (anonymous):

6?

jimthompson5910 (jim_thompson5910):

good, AC = 6

OpenStudy (anonymous):

Now what do I do?

jimthompson5910 (jim_thompson5910):

you just found AB, BC, and AC so you found all 3 sides of ABC, you're done

jimthompson5910 (jim_thompson5910):

AB = 6.67 BC = 5.33 AC = 6

OpenStudy (anonymous):

Oh so I dont have to add them all up?

jimthompson5910 (jim_thompson5910):

nope, they don't want the perimeter, just the side lengths of ABC

OpenStudy (anonymous):

Thank you soooo much for all your help tonight!!!!

jimthompson5910 (jim_thompson5910):

yw